### 29 Jan 2024 v1

# VIALM: A Survey and Benchmark of Visually Impaired Assistance with Large Models

Abstract
Visually Impaired Assistance (VIA) aims to automatically help visually impaired (VI) handle daily activities. The advancement of VIA primarily depends on developments in Computer Vision (CV) and Natural Language Processing (NLP), both of which exhibit cutting-edge paradigms with large models (LMs). Furthermore, LMs have shown exceptional multimodal abilities to tackle challenging physically-grounded tasks such as embodied robots. To investigate the potential and limitations of state-of-the-art (SOTA) LMs’ capabilities in VIA applications, we present an extensive study for the task of VIA with LMs (VIALM). In this task, given an image illustrating the physical environments and a linguistic request from a VI user, VIALM aims to output step-by-step guidance to assist the VI user in fulfilling the request grounded in the environment. The study consists of a survey reviewing recent LM research and benchmark experiments examining selected LMs’ capabilities in VIA. The results indicate that while LMs can augment VIA, their output cannot be well grounded in reality (i.e., 25.7% GPT-4’s responses) and lacks fine-grained guidance (i.e., 32.1% GPT-4’s responses).
1 Introduction
The term Visually Impaired (VI) refers to individuals with partial or complete blindness. It includes approximately 2.2 billion people worldwide. [^1] https://www.who.int/publications/i/item/9789241516570 To improve their lives, numerous previous studies have focused on Visually Impaired Assistance (VIA). It aims to provide support, tools, and services to aid the VI in their daily activities and promote their independence. The mainstream VIA research takes advantage of the advanced CV and NLP technology. For example, BrowseWithMe (Stangl et al., 2018) can talk with VI users via NLP to assist online shopping; ReCog (Ahmetovic et al., 2020) can help VI users understand the surrounding objects via CV.
To date, the predominant VIA efforts have been made with small models. However, large models (LMs) have recently transformed the landscape of CV, NLP, and multimodal research (Brown et al., 2020). For instance, in NLP, a Large Language Model (LLM) ChatGPT (Ouyang et al., 2022) showed superior cross-task language skills. Moreover, a large Visual-Language Model (VLM) GPT-4 (OpenAI, 2023) exhibited human-equivalent abilities in non-trivial multimodal tasks. More recently, a LLM-based embodied agent named PaLM-E (Driess et al., 2023) further grounded the vision and language capabilities to interact with physical environments.
Viewing this revolution, we raise a crucial question: can LMs transform the landscape of VIA? We consequently propose a task of Visual Impaired Assistance with Language Models (VIALM). The input of VIALM is an image illustrating the physical environment and a language request from a VI user about a task to complete in the environment. The output is a step-by-step environment-grounded guidance to assist the VI user in fulfilling their request. To illustrate, Figure 1 demonstrates a representative input and output of VIALM.
Figure 1: A sample input and output of VIALM. Its input is a pair of a visual image of the environment (the left image) and a user request in language (the grey box). The yellow box shows the output guidance for VI users to complete the request within the environment (the right image). The output should convey spatial information (blue words) and be comprehensible (green words) for VI users.
To the best of our knowledge, our study is the first to investigate LMs in VIA extensively. It consists of a thorough survey of important LMs’ work, including LLMs, VLMs, and embodied agents, all of which hold potential advancements for VIA. Subsequently, we present a fundamental benchmark for empirical studies by formulating the task following the Visual Question Answering (VQA) setup. It comprises 200 visual environment images with paired questions (user requests) and answers (guidance). The images cover two types of environments: supermarket and home. Based on the benchmark, we examined 6 SOTA VLMs, GPT-4, CogVLM (Wang et al., 2023b), Qwen-VL (Bai et al., 2023), LLaVA (Liu et al., 2023b), MiniGPT-v2 (Chen et al., 2023a), and BLIVE (Hu et al., 2024), to study their zero-shot capabilities in VIA.
The benchmark experiments reveal that the performance of current SOTA LMs is yet to be considered satisfying. We identify two main limitations in the LMs: (1) they struggle to generate environment-grounded guidance (25.7% of GPT-4’s responses), and (2) they lack tactile, fine-grained guidance (32.1% of GPT-4’s responses). Our experiments also indicate potential solutions: (1) addressing the former requires effective visual processing, while (2) the latter necessitates the fusion of tactile modality into language generation.
For future developments, we have open-sourced the resources [^2] https://github.com/YiyiyiZhao/VIALM. The contributions of this work are threefold:
$\bullet$ We introduce a novel task, VIALM, to extensively investigate how LMs transform the VIA landscape.
$\bullet$ We thoroughly survey the important LM work applicable to VIA and construct the first VIALM benchmark.
$\bullet$ We experiment with our benchmark and draw practical findings regarding SOTA LMs’ zero-shot VIA capabilities.
2 Large Models
Over the past two years, LMs have dramatically reshaped the fields of CV, NLP, and multimodal research. This section offers a comprehensive and current survey of LMs, focusing on a collection of 43 notable models. A chronological representation of the LMs is depicted in Figure 2. It is observed that the years 2021 and 2022 marked advancements in LLMs. This progress was subsequently followed by an upsurge in the development of VLMs and embodied agents in 2023.
2.1 Large Language Models
Name Release Time Pretraining Task Model Architecture # Maximum Parameters Instruction-tuned Variants
GPT-3 (Brown et al., 2020) 2020.05 175B
InstructGPT,
GPT-3.5
GLaM (Du et al., 2022) 2021.12 1.2T /
Gopher (Rae et al., 2021) 2021.12 280B /
LaMDA (Thoppilan et al., 2022) 2022.01 137B Bard (Manyika, 2023)
PaLM (Chowdhery et al., 2023) 2022.04 540B Flan-PaLM (Chung et al., 2022)
Chinchilla (Hoffmann et al., 2022) 2022.05 70B /
OPT (Zhang et al., 2022) 2022.05 175B OPT-IML (Iyer et al., 2022)
GLM-130B (Zeng et al., 2023) 2022.10 130B ChatGLM
BLOOM (Scao et al., 2022) 2022.11 176B BLOOMZ (Muennighoff et al., 2023)
Galactica (Taylor et al., 2022) 2022.11 120B Evol-Instruct
LLaMA (Touvron et al., 2023a) 2023.02 65B
Alpaca (Taori et al., 2023),
WizardLM,
Vicuna (Chiang et al., 2023)
RWKV (Peng et al., 2023) 2023.05 14B RWKV-4 Raven
PaLM-2 (Anil et al., 2023) 2023.05 / /
LLaMA-2 (Touvron et al., 2023b) 2023.07 70B
LLaMA2-Chat,
OpenChat V2
Falcon (Almazrouei et al., 2023) 2023.11 180B Falcon-instruct
Mixtral 8x7B (Jiang et al., 2024) 2024.01 13B Mixtral 8x7B–Instruct
Table 1: Summary of pretrained LLMs and their instruction-tuned variants, arranged chronologically from the earliest to the most recent releases. Categorized by pretraining tasks, symbols ➀, ➁, and ➂ denote language modeling (Radford et al., 2019), autoregressive blank infilling, and mixed objectives, respectively. Categorized by model architectures, symbols ➀, ➁, ➂ and ➃ represent transformer decoder (Radford et al., 2018), MOE decoder, bidirectional GLM, RWKV architecture, and encoder-decoder transformer, respectively. The symbol ”/” is used when specific information is not explicitly available. All these models have a maximum parameter count exceeding 10 billion. The majority of these LLMs were released in the past two years. It is evident that many instruction-tuned models are developed using LLaMA.
Language models are neural networks with text input and output, initially designed for NLP tasks. Large Language Models (LLMs) distinguish themselves by their extensive scale, indicated by the number of network parameters. An important feature of LLMs is their emergent abilities (Wei et al., 2022b), which have been leveraged in reasoning and decision-making (Yao et al., 2023). Following the definition of LLMs in Zhao et al. (2023a), we focus on models with more than 10 billion parameters. In addition, this survey categorizes LLMs based on architectures (Brown et al., 2020) and pretraining tasks (Zeng et al., 2023) and investigates instruction-tuned LLMs (Wei et al., 2022a), systematically summarized in Table 1.
Pretraining Task
The most widely adopted LLM pretraining task is language modeling (Radford et al., 2019), which predicts the next tokens in an autoregressive manner. The success of GPT series (Brown et al., 2020) has demonstrated this approach’s effectiveness, especially when the language model scale is significantly increased and the training corpus is extensively expanded. Consequently, a substantial number of LLMs have adopted it, including GLaM (Du et al., 2022), Gopher (Rae et al., 2021), LaMDA (Thoppilan et al., 2022), PaLM (Chowdhery et al., 2023), Chinchilla (Hoffmann et al., 2022), OPT (Zhang et al., 2022), BLOOM (Scao et al., 2022), Galactica (Taylor et al., 2022), LLaMA (Touvron et al., 2023a), RWKV (Peng et al., 2023), LLaMA-2 (Touvron et al., 2023b), Falcon (Almazrouei et al., 2023) and Mixtral (Jiang et al., 2024). Meanwhile, some LLMs have implemented mixed schemes to enhance language understanding. For example, GLM-130B has introduced autoregressive blank infilling (Zeng et al., 2023), and more recently, PaLM-2 has adopted a mixture of objectives (Anil et al., 2023).
Model Architecture
The model architectures generally align with pretraining tasks. Predominantly, LLMs utilize the transformer decoder architecture, which synergizes well with language modeling pretraining. Some models have embraced this architecture, including GPT-3, Gopher, LaMDA, PaLM, Chinchilla, OPT, BLOOM, Galactica, LLaMA, LLaMA-2, and Falcon. Additionally, there are variations in model architecture. For instance, GLaM and Mixtral employ the Mixture of Experts (MOE) approach, GLM-130B features a bidirectional design, RWKV introduces their unique architecture, and PaLM-2 has adopted the encoder-decoder architecture.
Instruction-tuned LLMs
Instruction tuning employs a limited instruction dataset combined with multitask fine-tuning, showing advantages in aligning LLMs with human intentions (Wei et al., 2022a; Li et al., 2023c). Many fundamental LLMs have related instruction-tuned variants. One category of variants, with ChatGPT being the most renowned, is those capable of functioning in a conversational manner. Additionally, LLaMA and LLaMA-2 have become prominent fundamental models for creating instruction-tuned LLMs. Notable tuned examples include Alpaca (Taori et al., 2023), Vicuna (Chiang et al., 2023), LLaMA2-Chat, among others.
Figure 2: Timeline of LMs. Release times are based on the publication dates of their respective descriptive papers on arXiv. LLMs are marked in black, large VLMs in blue, and embodied agents in green. It is observed the years 2021 and 2022 saw LLM advancements, followed by a 2023 surge in VLMs and embodied agent development.
2.2 Large Vision-Language Models
Large Vision-Language Models (VLMs) process both visual and language inputs, enabling them to handle multimodal tasks such as image captioning and VQA. In the context of VIA, we focus on VLMs that generate linguistic outputs, ensuring accessibility for VI users. A typical VLM architecture consists of three key components: a visual encoder, a LLM, and a vision-language connector (Liu et al., 2023b). This survey categorizes VLMs according to their model components and training/tuning strategies, as summarized in Table 2.
Name Release Time Visual Encoder LLM Connector Training/Tuning
Flamingo
(Alayrac et al., 2022)
2022.04
NFNet
(Brock et al., 2021)
Chinchilla Cross-attention dense layers Pretraining
PaLI
(Chen et al., 2023d)
2022.09
ViT-e
(Zhai et al., 2022)
mT5 (Raffel et al., 2020) / Multi-task Pretraining
BLIP-2
(Li et al., 2023a)
2023.01
1. ViT-L/14
(Radford et al., 2021),
2. ViT-g/14
(Fang et al., 2023)
1. OPT
2. FlanT5
Q-former Multi-task Pretraining
GPT-4
(OpenAI, 2023)
2023.03 / GPT-4 / /
LLaVa
(Liu et al., 2023b)
2023.04 ViT-L/14 LLaMA Projection layer
1. Pretraining,
2. Instruction tuning.
PaLI-X
(Chen et al., 2023c)
2023.05
ViT-22B
(Dehghani et al., 2023)
UL2 Projection layer
1. Multi-task pretraining,
2. Task-specific fine-tuning.
InstructBLIP
(Dai et al., 2023)
2023.05 ViT-g/14
1. FlanT5
2. Vicuna
Q-former
1. Pretraining,
2. Instruction tuning.
VisionLLM
(Wang et al., 2023c)
2023.05
1. ResNet,
2. InternImage-H
Alpaca
BERT-Base and
Deformable DETR (Zhu et al., 2021)
Multi-task pretraining
BLIVA
(Hu et al., 2024)
2023.08 ViT-g/14 FlanT5 Q-former and projection layer
1. Pretraining,
2. Instruction-tuning.
Qwen-VL
(Bai et al., 2023)
2023.08
ViT
(Dosovitskiy et al., 2021)
Qwen Single-layer cross-attention module
1. Pretraining,
2. Multi-task training,
3. Conversation tuning.
MiniGPT-v2
(Chen et al., 2023a)
2023.10 ViT-g/14 LLaMA2-Chat Linear projection layer
1. Pretraining,
2. Multi-task training,
3. Instruction tuning.
CogVLM
(Wang et al., 2023b)
2023.11
Eva-clip ViT
(Sun et al., 2023)
Vicuna MLP adapter
1. Pretraining;
2. Multi-task training
Gemini
(OpenAI, 2023)
2023.12 / PaLM / /
Table 2: Summary of pretrained VLMs. A large VLM typically comprises a visual encoder, a LLM, and a connector. It is indicated that ViT is the most favored visual encoder, while LLMs primarily derive from the LLaMA family. The most common connectors are projection layers. Pretraining for VLMs generally encompasses three phases: foundational training, multi-task fine-tuning, and instructional tuning.
Visual Encoder
Table 2 indicates that ViT is the most preferred visual encoder in current VLMs. Specifically, Qwen-VL implements the standard ViT framework, whereas PaLI-X utilizes the highly scaled-up 22B variant. The ViT-L/14 and ViT-g/14 models, both examined in the BLIP-2 study, are particularly popular. LLaVA has adopted ViT-L/14, while InstructBLIP, BLIVA, and MiniGPT-v2 have selected ViT-g/14. The ViT transformer architecture offers substantial benefits in visual encoding, including scalability in both model and data size (Dehghani et al., 2023), and a more effective integration of visual and language embeddings (Fang et al., 2023).
LLM
In the early stage of the development of VLMs, several models, including BLIP-2, InstructBLIP, and BLIVA, chose the instruction-tuned pioneer FlanT5 as their foundational LLM component. Recently, the LLaMA family, known for its open-source accessibility and impressive performance in NLP tasks, has attracted significant attention. The release of LLaMA-2 is expected to continue and possibly foster this trend. Numerous VLMs, including LLaVA, MiniGPT-4, InstructBLIP, MiniGPT-v2, and CogVLM, have been developed using language models from the LLaMA family, such as Vicuna, Alpava, and LLaMA-2-Chat, as shown in Table 2.
Connector
BLIP-2 introduced the Q-Former component, which was subsequently adopted by InstructBLIP and BLIVA. However, the most common design for connectors currently is a linear projection layer or a multi-layer perceptron(MLP). Models like LLaVA, PaLI-X, MiniGPT-v2, and CogVLM have adopted this design. Similarly, Qwen-VL has adopted the linear layer connection approach, but it improves its functionality by adding cross-attention mechanisms. Comparing MLP with Q-Former, the LLaVA-v1.5 study (Liu et al., 2023a) suggests that the Q-Former may have shortcomings in effectively managing the output length of LLMs.
Training/Tuning
The prevalent pre-training methodology for VLMs generally consists of three phases. Initially, the fundamental pretraining phase utilizes extensive image-text pairs from the Internet to equip the model with general capabilities for visual-language tasks. Subsequently, models undergo multi-task fine-tuning, during which the dataset selection is more carefully curated. The final stage is centered on instruction tuning. It utilizes a limited amount of instructional or conversational data to optimize the model for interactive user engagement. During these stages, the training components of the models may differ. In the initial stage, the connector may update its parameters. However, models typically do not update the LLM parameters at this stage due to their substantial magnitude. In contrast, many VLMs update the LLM parameters during the final instruction tuning phase. Models with frozen LLM include MiniGPT-4 and BLIVA. Although CogVLM does not update its own LLM parameters, it incorporates an additional module into the LLM and updates this module’s parameters. Models that update LLM parameters include LLaVA, Qwen-VL, and MiniGPT-v2.
2.3 Embodied Agents
LLM-based agents, especially embodied agents, have demonstrated rapid and significant emergence across various scenarios (Xi et al., 2023). The specialized embodied components enable these agents to interact with environments during problem-solving by perceiving the surroundings and performing actions. These actions can either be autonomously carried out by the agent itself or with the aid of additional tools. In this study, we focus on embodied agents that perform actions independently, rather than relying on external tools. This survey analyzes the use of LLMs/VLMs, categorizing their tasks, environments, and design strategies for transforming them into autonomous agents proficient in task-solving. A systematic summary can be found in Table 3.
Name Time LLM Task Environment LM Output Code Interface LM Prompting/Tuning
LM as Planner (Huang et al., 2022a) 2022.01 GPT3, CodeX VirtualHome tasks
SayCan (Ichter et al., 2022) 2022.04 PaLM Real-world robotic tasks
Code as Policy (Liang et al., 2023) 2022.09 Codex Real-world robotic tasks
Inner Monologue (Huang et al., 2022b) 2022.07 InstructGPT Ravens tasks and real-world robotic tasks ➀ + ➁
ProgPrompt (Singh et al., 2023) 2022.09 GPT-3 VituralHome tasks and real-world robotic tasks ➀ + ➁
LLM-Planner (Song et al., 2022) 2022.12 GPT-3 Alfred tasks
Matcha (Zhao et al., 2023b) 2023.03 GPT-3 CoppeliaSim-simulated NICOL robot tasks
PaLM-E (Driess et al., 2023) 2023.03 PaLM TAMP tasks and real-world robotic tasks
Grounded Decoding (Huang et al., 2023) 2023.03 InstructGPT, PaLM Ravens tasks and real-world robotic tasks ➀ + ➁
Voyager (Wang et al., 2023a) 2023.05 GPT-4 Minecraft Tasks
GITM (Zhu et al., 2023) 2023.05 GPT-3.5 Minecraft Tasks
RT-2 (Brohan et al., 2023) 2023.07 PaLI-X,PaLM-E Real-world robotic tasks
ViLA (Hu et al., 2023) 2023.11 GPT-4V Ravens tasks and real-world robotic tasks ➀ + ➁
AutoRT (Ahn et al., 2024) 2024.01 / Real-world robotic tasks
Table 3: Summary of LLM-based embodied agents. Categorized by environment, ➀ symbolizes simulated enrionments and ➁ the physical world. In LM output, ➀ means mid-level plans, while ➁ indicates low-level robotic actions. Regarding code interface, ➀ is code-free (language-only), and ➁ involves programming code. In LM operations, ➀ is for prompting, ➁ for tuning. It is observed that most embodied agents employ LM prompting for mid-level plans in natural language, which are then translated into low-level robotic actions.
LLM
GPT-3.5 and GPT-4, accessible through API services, are the favored LLMs for developing embodied agents. As shown in Table 3, models such as Inner Monologue (Huang et al., 2022b), ProgPrompt (Singh et al., 2023), LLM-Planner (Song et al., 2022), Matchat (Zhao et al., 2023b), Grounded Decoding (Huang et al., 2023), Voyager (Wang et al., 2023a), GITM (Zhu et al., 2023), and ViLA (Hu et al., 2023) utilize GPT models as the cognitive core for agents’ thinking and planning processes. Additionally, ’Code as Policy’ (Liang et al., 2023) employs CodeX, a derivative of the GPT series focusing on programming. Another prominent LLM is PaLM, which has been chosen by agents like SayCan (Ichter et al., 2022), Grounded Decoding, and RT-2 (Brohan et al., 2023).
Environment and Task
There are two types of environments in which embodied agents operate: simulated and real-world environments. Within simulated settings, agents are trained to execute predefined skills. For instance, Voyager and GITM are designed for playing Minecraft. Additionally, some agents, functioning as virtual robots, utilize platforms like VirtualHome (Puig et al., 2018) for cost-effective training prior to potential deployment in real-world scenarios. Furthermore, numerious embodied agents are able to execute in the physical world. These include SayCan, ’Code as Policy’, Inner Monologue, ProgPrompt, Matcha, PaLM-E, Grounded Decoding, RT-2, ViLA, AutoRT (Ahn et al., 2024), and so on.
From LLM to Agent
To transform LLMs into agents, a typical method is a two-tier framework: the LLM initially generates natural language sequences as high-level plans, later translated into executable low-level actions. Agents using this approach include Inner Monologue, LLM-Planner, Matcha, PaLM-E, Grounded Decoding, ViLA and AutoRT. Alternatively, some models have implemented an end-to-end action generation scheme. Such agents include SayCan, ’Code as Policy’, ProgPrompt, Voyager, GITM, and RT-2. Notably, RT-2 treats action as a modality and has introduced the concept of a visual-language-action model. In addition, while many agents leverage the prompt approach in utilizing LLMs, RT-2 fine-tunes the LLMs for improved performance. Moreover, as next-step action generation is inherently logical, some research conveys a preference for using code instead of natural language to facilitate this process. Agents that have implemented programming code for action generation include ProgPrompt, ’Code as Policy’, and Voyager.
3 Benchmark Evaluation
In order to explore the question ’Can LMs transform the landscape of VIA?’ and to develop a roadmap for VIALM, we conducted a benchmark evaluation. This assessment involves measuring and comparing the zero-shot performance of LMs, aiming to help readers understand the models’ strengths and weaknesses in the following four key aspects. (1) Viability: Do LMs possess the potential and capabilities necessary for VIA? (2) Environment Adaptability: If LMs are effective for VIA, can they maintain their effectiveness in complex environments? (3) Functionality: Can LMs achieve effective performance in real-world practice? (4) Model Selection: How do different designs of LMs affect performance in VIA?
Evaluation Models
Chen et al. (2023b) demonstrated that end-to-end VLMs outperform separate visual and language pipelines in embodied decision-making tasks. We selected six SOTA end-to-end VLMs for our evaluation, namely GPT-4, CogVLM, MiniGPT, Qwen-VL, LLaVA and BLIVA. The specific versions used were GPT-4, CogVLM-17B-Chat-V1.1, MiniGPT-v2 with LLaMA-2-Chat-7B, Qwen-VL-Chat with Qwen-7B, LLaVA-v1.5-7B, and Bliva-vicuna-7B. All models operated with float16 precision, with the exception of GPT-4.
3.1 Evaluation Dataset
Given that VIALM processes environmental information and user requests to ultimately generate step-by-step guidance, the benchmark is structured to follow the VQA framework. The input comprises an image depicting the scene in front of the VI user, combined with a question reflecting their intention. The output is an answer that delivers the desired guidance. Due to the scarcity of datasets specifically tailored for the VIALM task and the widespread use of popular VQA datasets (Marino et al., 2019; Gurari et al., 2018) in the fine-tuning and instruction tuning of VLMs (Dai et al., 2023), we have carefully developed an evaluation dataset. This dataset is designed to (1) accurately reflect, measure, and compare performance on the VIALM task, and (2) remain unexposed to these VLMs during their training and tuning phases.
Dataset Annotation
Figure 3: Evaluation Data Sample.The test dataset is in the format of a VQA dataset. The question indicates that the VI user is searching for a toilet. The output answer offers detailed guidance related to this environment to assist the user in achieving this objective.
We collected images from the Internet of the two most common daily life environments: indoor homes and supermarkets. The evaluation dataset was subsequently manually crafted by employing highly educated annotators with English-medium undergraduate or graduate degrees (Li et al., 2023b). Figure 3 illustrates a test data sample.
Dataset Statistics
The evaluation dataset consists of 200 samples in total, divided into two scenarios with 100 images each, similar in size to Chen et al. (2023b). Every image is paired with a corresponding question and answer. The questions are designed to prompt the models to offer guidance for VI users, utilizing tactile clues to navigate towards specific targets. Notably, the dataset features a broad coverage, with 96 different target objects, such as ’meat’ and ’basin’ in the supermarket scenario, and ’sofa’ and ’fridge’ in the home scenario. Furthermore, in comparison to benchmark datasets like OK-VQA (average answer length: 8.1 tokens) (Marino et al., 2019) and CLEVR (average answer length: 18.4 tokens) (Johnson et al., 2017), our dataset exhibits longer questions (average question length: 51 tokens) and significantly longer answers (average answer length: 99.8 tokens). It indicates a higher level of difficulty for models in question answering.
3.2 Evaluation Metrics
We utilize a hierarchical set of metrics across various dimensions for evaluation. These metrics fall into two main categories: Automatic Evaluation and Human Evaluation.
Automatic Evaluation Metrics
We analyze the generated responses (Ni’mah et al., 2023) using both surface-level and semantic-level metrics. Surface-level metrics evaluate the overlap of tokens or strings between generated and reference texts. For this, we adopt the ROUGE metric (Lin, 2004), which is the standard for document summarization. Semantic-level metrics assess the semantic coherence and relevance of the model’s outputs in comparison to the ground truths. For this, we use BERTScore (Zhang et al., 2020), a widely recognized metric that evaluates semantic similarity.
Human Evaluation Metrics
The human evaluation measures the models’ performance in three main aspects: Correctness, Actionability, and Clarity, culminating in an overall helpfulness score. Correctness evaluates the accuracy of the guidance in describing the environment and indicating target positions. Actionability evaluates the user’s ease in following the provided guidance. Clarity assesses the comprehensibility of the answer’s language from the user’s perspective. The overall helpfulness score is calculated using weights of 0.5 for Actionability, 0.4 for Correctness, and 0.1 for Clarity.
4 Experiment Results
Figure 4: Evaluation Results. (a) Overall average experimental results for six models—GPT-4, CogVLM, MiniGPT, Qwen-VL, LLaVA, BLIVA—based on ROUGE, BERTScore, and Human Evaluation metrics. (b) Human evaluation results of the six models focusing on the aspects of Correctness, Actionability, and Clarity.
We present the overall experimental results in Figure 4 (a). First, GPT-4 and LLaVA both exhibit relatively high BERTSCore values (above 0.6 within a range of 0 to 1), indicating that current SOTA VLMs possess the potential for the VIALM task (Viability). Second, regarding ROUGE, all models demonstrate similar but low values (below 0.2 within a range of 0 to 1), suggesting that the model predictions may be further improved. Third, the human evaluation results, which have been scaled into a range from 0 to 1, reveal that GPT-4 performs the best, with LLaVA and CogVLM following as the second-best performers, motivating further study.
4.1 Automatic Evaluation Results
Supermarket
Model RG_1 RG_2 RG_L BERT_S Len
GPT-4 0.246 0.047 0.138 0.614 229.39
CogVLM 0.319 0.063 0.187 0.477 65.22
MiniGPT4 0.269 0.046 0.15 0.514 115.71
Qwen-VL 0.254 0.05 0.156 0.505 151.31
LLaVA 0.283 0.053 0.168 0.592 139.74
BLIVA 0.208 0.03 0.123 0.494 192.07
Home
Model RG_1 RG_2 RG_L BERT_S Len
GPT-4 0.356 0.082 0.183 0.698 195.07
CogVLM 0.324 0.085 0.174 0.521 52.7
MiniGPT4 0.331 0.081 0.182 0.597 97.85
Qwen-VL 0.357 0.087 0.19 0.598 88.13
LLaVA 0.374 0.101 0.201 0.662 103.78
BLIVA 0.317 0.073 0.175 0.639 195.79
Table 4: Automatic Evaluation Results. Specific versions are GPT-4, CogVLM-17B-Chat-V1.1, MiniGPT-v2-7B, Qwen-VL-Chat-7B, LLaVA-v1.5-7B, and Bliva-vicuna-7B. Except for GPT-4, the others operated with float16 precision. The highest and runner-up metric values are distinctly marked in gray. ’RG’ represents ROUGE, ’BERT_S’ stands for BERTScore, and ’Len’ signifies output length.
Table 4 displays automatic evaluation results for the two distinct scenarios. In both situations,GPT-4 and LLaVA possess the highest BERTScores,indicating their superior abilities. Additionally, when considering ROUGE, we can see: (1) all models perform better in the home scenario setting; (2) LLaVA displays relatively high ROUGE values and CogVLM excels in a more challenging environment; (3) GPT-4 performs worse than smaller-scaled models. The first observation suggests that the increased visual complexity of the supermarket environment (i.e., a greater variety of objects and more colors in the image) makes the VIALM task more challenging, highlighting the need for models to generalize with environment adaptability. Combining the second and third observations and analyzing the generated guidance, we identify GPT-4’s significant drawback: its output is excessively redundant and lengthy, resulting in low ROUGE scores and reduced functionality. In contrast, CogVLM generates the shortest outputs, while LLaVA exhibits the best performance with a moderate guidance length and high ROUGE scores.
4.2 Human Evaluation Results
Model Not Grounded Rate Not Fine-grained Rate
GPT-4 0.257 0.321
CogVLM 0.45 0.721
MiniGPT4 0.736 0.8
Qwen-VL 0.607 0.743
LLaVA 0.486 0.543
BLIVA 0.85 0.886
Table 5: Human Evaluation Fail Rates. ’Not Grounded Rate’ is calculated by the sample proportion with Correctness $\leq$ 3 (Score range: 1–5). ’Not Fine-grained Rate’ is derived from Actionability $\leq$ 3.
Figure 4 (b) presents the human evaluation results. To mitigate bias, two independent participants conducted evaluations. In each distinct scenario, 50 random samples were selected. Pearson and Spearman correlation coefficients between the two annotators for each dimension range from 0.7 to 0.9, indicating a strong consensus on the models’ performance. First, it is once again observed that all models provide more accurate answers in home settings than in supermarket settings. Second, as shown in Figure 4 (b), GPT-4 excels in performance across all three dimensions, with CogVLM and LLaVA closely following in second place. Subsequently, we analysis the results for the three specific dimensions:
$\bullet$ In Correctness, GPT-4 performs the best, with CogVLM slightly surpassing LLaVA. It might indicate that CogVLM understands the visual environment better than LLaVA, and therefore leads to a superior environment grounding ability.
$\bullet$ In Actionability, GPT-4 performs the best, with LLaVA narrowly outperforming CogVLM. Since we prompted the use of tactile clues, this observation implies LLaVA’s LLM component, LLaMA-2, generates more effective guidance to VI users, incorporating tactile and fine-grained instructions.
$\bullet$ In Clarity, GPT-4 obtains the highest score. Nevertheless, its responses tend to be redundant and long-winded, necessitating extra effort to achieve functionality. In contrast, other models provide responses that are not excessively long but tend to be overly simplistic, lacking sufficient information.
This analysis reveals two prevalent shortcomings shared by all six models, summarized as ’not grounded’(indicating low correctness) and ’not fine-grained’ (indicating low actionability). The failure rates are presented in Table 5. It is notable that, despite GPT-4’s superior performance, 25.7% of its responses are not grounded and 32.1% not fine-grained.
4.3 Case Study
To demonstrate the models’ output, we present an example case in Figure 5 where the question seeks guidance to obtain mustard. The predictions of GPT-4, CogVLM, and LLaVA are provided, yielding the following observations: (1) GPT-4 generates numerous tokens offering general guidance but lacks image grounding. In addition, its guidance is not easy for VI users to follow. (2) CogVLM exhibits better environment grounding, specifying the object’s position to the user’s left, yet it lacks detailed step-by-step guidance. (3) LLaVA’s output emphasizes guiding the user to the specific object but appears to have limited understanding of the environment.
Figure 5: An example of predictions from the top three models. GPT-4 is prone to generating redundant output describing general situations. CogVLM exhibits advantages in image grounding, while LLaVA emphasizes providing step-by-step guidance for VI users.
4.4 Discussions
Home
Model Rouge_1 Rouge_2 Rouge_L BERTScore
Grounded Part
GPT-4 0.156 0.028 0.104 0.405
CogVLM 0.262 0.043 0.158 0.284
MiniGPT4 0.211 0.053 0.139 0.415
Qwen-VL 0.201 0.028 0.132 0.311
LLaVA 0.186 0.039 0.123 0.404
BLIVA 0.133 0.025 0.091 0.375
Fine-Grained Part
GPT-4 0.292 0.072 0.166 0.573
CogVLM 0.317 0.083 0.189 0.443
MiniGPT4 0.295 0.062 0.177 0.455
Qwen-VL 0.331 0.088 0.193 0.506
LLaVA 0.341 0.094 0.202 0.544
BLIVA 0.259 0.062 0.159 0.514
Table 6: Automatic Evaluation Results with separated guidance parts extracted from ground truths. Metric scores are calculated by comparing model predictions with each corresponding part.
This section explores model selection strategies to improve VIALM, particularly in addressing the two mentioned limitations. We divided the ground truth text into two parts: the grounded part (i.e., environmental), and the fine-grained part (i.e., tactile). We then compared model predictions with each part and calculated automatic metric scores. Since Table 4 reveals that Qwen-VL and LLaVA exhibit superior performance under the home scenario, a notable deviation from the supermarket scenario and our human evaluation results, we focus on the home scenario. Results are displayed in Table 6.
Analyzing Table 6 reveals two significant findings. Firstly, CogVLM excels in environment grounding, possibly due to two factors: (1) the incorporation of the more recent Eva-Clip ViT model, enhancing its visual perception capabilities, and (2) the inclusion of a Visual Expert module in CogVLM’s LLM architecture, augmenting its visual-language alignment abilities. Secondly, LLaVA demonstrates superior efficiency in generating fine-grained guidance, possibly because its LLM component, LLaMA-2, might possess advanced capabilities concerning tactile description. These initial findings can assist researchers in developing VIALM models.
5 Conclusion
In this study, we defined the VIALM task and carried out a comprehensive survey of LMs for VIA, encompassing LLMS, VLMs, and embodied agents. Furthermore, we constructed a benchmark dataset with comprehensive coverage and high difficulty. Subsequently, we conducted a benchmark evaluation to measure the zero-shot performance of LMs, aiming to examine the potential and limitations of current SOTA models in the VIALM task. The experimental results revealed two primary limitations in the models: their ability to generate grounded and fine-grained guidance. To address these limitations, the experiment results suggest that improving visual processing can aid in producing more grounded advice, and incorporating the tactile modality can contribute to generating more fine-grained guidance for VI users.

### 10 Feb 2024 v2

@@
 Abstract
-Visually Impaired Assistance (VIA) aims to automatically help visually impaired (VI) handle daily activities.
+Visually Impaired Assistance (VIA) aims to automatically help the visually impaired (VI) handle daily activities.
 The advancement of VIA primarily depends on developments in Computer Vision (CV) and Natural Language Processing (NLP), both of which exhibit cutting-edge paradigms with large models (LMs).
@@
 The study consists of a survey reviewing recent LM research and benchmark experiments examining selected LMs’ capabilities in VIA.
-The results indicate that while LMs can augment VIA, their output cannot be well grounded in reality (i.e., 25.7% GPT-4’s responses) and lacks fine-grained guidance (i.e., 32.1% GPT-4’s responses).
+The results indicate that while LMs can potentially benefit VIA, their output cannot be well environment-grounded (i.e., 25.7% GPT-4’s responses) and lacks fine-grained guidance (i.e., 32.1% GPT-4’s responses).
 1 Introduction
@@
 More recently, a LLM-based embodied agent named PaLM-E (Driess et al., 2023) further grounded the vision and language capabilities to interact with physical environments.
+Figure 1: A sample input and output of VIALM.
+Its input is a pair of a visual image of the environment (the left image) and a user request in language (the grey box).
+The yellow box shows the output guidance for VI users to complete the request within the environment (the right image).
+The output should convey environment-grounded information (blue words) and fine-grained (green words) guidance, integrating tactile support (the last sentence) for VI users.
 Viewing this revolution, we raise a crucial question: can LMs transform the landscape of VIA?
@@
 The output is a step-by-step environment-grounded guidance to assist the VI user in fulfilling their request.
-To illustrate, Figure 1 demonstrates a representative input and output of VIALM.
-Figure 1: A sample input and output of VIALM.
-Its input is a pair of a visual image of the environment (the left image) and a user request in language (the grey box).
-The yellow box shows the output guidance for VI users to complete the request within the environment (the right image).
-The output should convey spatial information (blue words) and be comprehensible (green words) for VI users.
+This guidance need to be easy to follow (fine-grained) for these users.
+Particularly, it is crucial to involve tactile guidance as VI users heavily rely on tactile cognition to complement vision loss (Ottink et al., 2022).
+Figure 1 shows a sample input and output of VIALM.
 To the best of our knowledge, our study is the first to investigate LMs in VIA extensively.
-It consists of a thorough survey of important LMs’ work, including LLMs, VLMs, and embodied agents, all of which hold potential advancements for VIA.
+It consists of a thorough survey of important LMs’ work, including LLMs, VLMs, and embodied agents, all of which hold potential benefits for VIA.
 Subsequently, we present a fundamental benchmark for empirical studies by formulating the task following the Visual Question Answering (VQA) setup.
-It comprises 200 visual environment images with paired questions (user requests) and answers (guidance).
+It comprises 200 visual environment images with paired questions (user requests) and answers (VIA guidance).
 The images cover two types of environments: supermarket and home.
 Based on the benchmark, we examined 6 SOTA VLMs, GPT-4, CogVLM (Wang et al., 2023b), Qwen-VL (Bai et al., 2023), LLaVA (Liu et al., 2023b), MiniGPT-v2 (Chen et al., 2023a), and BLIVE (Hu et al., 2024), to study their zero-shot capabilities in VIA.
-The benchmark experiments reveal that the performance of current SOTA LMs is yet to be considered satisfying.
-We identify two main limitations in the LMs: (1) they struggle to generate environment-grounded guidance (25.7% of GPT-4’s responses), and (2) they lack tactile, fine-grained guidance (32.1% of GPT-4’s responses).
-Our experiments also indicate potential solutions: (1) addressing the former requires effective visual processing, while (2) the latter necessitates the fusion of tactile modality into language generation.
-For future developments, we have open-sourced the resources [^2] https://github.com/YiyiyiZhao/VIALM. The contributions of this work are threefold:
+Benchmark experiments identify two main limitations in current SOTA LMs: (1) inferiority to generate environment-grounded guidance (25.7% of GPT-4’s responses), and (2) a lack of fine-grained guidance (32.1% of GPT-4’s responses), with a shortfall in integrating tactile sensation.
+However, it is also revealed that visually-focused LMs excel in environment understanding, while those with advanced language models are superior in generating easy-to-follow guidance.
+To overcome these limitations, we propose potential solutions: (1) improve visual capabilities to enhance environmental grounding, and (2) incorporate tactile modalities into language generation to produce more fine-grained guidance.
+A promising direction for future research involves synergistically enhancing multimodal capabilities to boost overall effectiveness.
+For future developments, we have open-sourced the resources. [^2] https://github.com/YiyiyiZhao/VIALM The contributions of this work are threefold:
 $\bullet$ We introduce a novel task, VIALM, to extensively investigate how LMs transform the VIA landscape.
@@
 Over the past two years, LMs have dramatically reshaped the fields of CV, NLP, and multimodal research.
-This section offers a comprehensive and current survey of LMs, focusing on a collection of 43 notable models.
-A chronological representation of the LMs is depicted in Figure 2.
-It is observed that the years 2021 and 2022 marked advancements in LLMs.
-This progress was subsequently followed by an upsurge in the development of VLMs and embodied agents in 2023.
+This section offers a comprehensive and current survey of LMs, focusing on 43 notable examples, chronologically depicted in Figure 2.
+It is observed that the years 2021 and 2022 marked advancements in LLMs, followed by an upsurge in the development of VLMs and embodied agents in 2023 and 2024.
+Figure 2: Timeline of LMs.
+Release times are based on the publication dates of their respective descriptive papers on arXiv.
+LLMs are marked in black, large VLMs in blue, and embodied agents in green.
+It is observed the years 2021 and 2022 saw LLM advancements, followed by a 2023 surge in VLMs and embodied agent development.
 2.1 Large Language Models
@@
 Language models are neural networks with text input and output, initially designed for NLP tasks.
-Large Language Models (LLMs) distinguish themselves by their extensive scale, indicated by the number of network parameters.
-An important feature of LLMs is their emergent abilities (Wei et al., 2022b), which have been leveraged in reasoning and decision-making (Yao et al., 2023).
-Following the definition of LLMs in Zhao et al.
-(2023a), we focus on models with more than 10 billion parameters.
-In addition, this survey categorizes LLMs based on architectures (Brown et al., 2020) and pretraining tasks (Zeng et al., 2023) and investigates instruction-tuned LLMs (Wei et al., 2022a), systematically summarized in Table 1.
-Pretraining Task
-The most widely adopted LLM pretraining task is language modeling (Radford et al., 2019), which predicts the next tokens in an autoregressive manner.
+Large Language Models (LLMs) stand out by their extensive scale, as measured by network parameters.
+A key feature of LLMs is their emergent abilities (Wei et al., 2022b), which have been leveraged in reasoning and decision-making (Yao et al., 2023).
+We focus on models with more than 10 billion parameters, in line with the definition in Zhao et al.
+(2023a).
+This survey categorizes LLMs based on architectures and pretraining tasks and investigates instruction-tuned LLMs, summarized in Table 1.
+Pretraining Tasks.
+The predominant pretraining task for LLMs is language modeling (Radford et al., 2019), which predicts the next tokens in an autoregressive manner.
 The success of GPT series (Brown et al., 2020) has demonstrated this approach’s effectiveness, especially when the language model scale is significantly increased and the training corpus is extensively expanded.
-Consequently, a substantial number of LLMs have adopted it, including GLaM (Du et al., 2022), Gopher (Rae et al., 2021), LaMDA (Thoppilan et al., 2022), PaLM (Chowdhery et al., 2023), Chinchilla (Hoffmann et al., 2022), OPT (Zhang et al., 2022), BLOOM (Scao et al., 2022), Galactica (Taylor et al., 2022), LLaMA (Touvron et al., 2023a), RWKV (Peng et al., 2023), LLaMA-2 (Touvron et al., 2023b), Falcon (Almazrouei et al., 2023) and Mixtral (Jiang et al., 2024).
+Consequently, a substantial number of LLMs have adopted it, including GLaM (Du et al., 2022), Gopher (Rae et al., 2021), LaMDA (Thoppilan et al., 2022), PaLM (Chowdhery et al., 2023), Chinchilla (Hoffmann et al., 2022), OPT (Zhang et al., 2022), BLOOM (Scao et al., 2022), Galactica (Taylor et al., 2022), LLaMA (Touvron et al., 2023a), RWKV (Peng et al., 2023), LLaMA-2 (Touvron et al., 2023b), Falcon (Almazrouei et al., 2023), and Mixtral (Jiang et al., 2024).
 Meanwhile, some LLMs have implemented mixed schemes to enhance language understanding.
 For example, GLM-130B has introduced autoregressive blank infilling (Zeng et al., 2023), and more recently, PaLM-2 has adopted a mixture of objectives (Anil et al., 2023).
-Model Architecture
+Model Architectures.
 The model architectures generally align with pretraining tasks.
@@
 For instance, GLaM and Mixtral employ the Mixture of Experts (MOE) approach, GLM-130B features a bidirectional design, RWKV introduces their unique architecture, and PaLM-2 has adopted the encoder-decoder architecture.
-Instruction-tuned LLMs
+Instruction-tuned LLMs.
 Instruction tuning employs a limited instruction dataset combined with multitask fine-tuning, showing advantages in aligning LLMs with human intentions (Wei et al., 2022a; Li et al., 2023c).
@@
 Notable tuned examples include Alpaca (Taori et al., 2023), Vicuna (Chiang et al., 2023), LLaMA2-Chat, among others.
-Figure 2: Timeline of LMs.
-Release times are based on the publication dates of their respective descriptive papers on arXiv.
-LLMs are marked in black, large VLMs in blue, and embodied agents in green.
-It is observed the years 2021 and 2022 saw LLM advancements, followed by a 2023 surge in VLMs and embodied agent development.
 2.2 Large Vision-Language Models
@@
 (Alayrac et al., 2022)
-2022.04
-NFNet
-(Brock et al., 2021)
-Chinchilla Cross-attention dense layers Pretraining
+2022.04 NFNet Chinchilla Cross-attention dense layers Pretraining
 PaLI
@@
 Pretraining for VLMs generally encompasses three phases: foundational training, multi-task fine-tuning, and instructional tuning.
-Visual Encoder
+Visual Components.
 Table 2 indicates that ViT is the most preferred visual encoder in current VLMs.
-Specifically, Qwen-VL implements the standard ViT framework, whereas PaLI-X utilizes the highly scaled-up 22B variant.
+Qwen-VL implements the standard ViT framework, whereas PaLI-X utilizes the highly scaled-up 22B variant.
 The ViT-L/14 and ViT-g/14 models, both examined in the BLIP-2 study, are particularly popular.
@@
 The ViT transformer architecture offers substantial benefits in visual encoding, including scalability in both model and data size (Dehghani et al., 2023), and a more effective integration of visual and language embeddings (Fang et al., 2023).
-LLM
-In the early stage of the development of VLMs, several models, including BLIP-2, InstructBLIP, and BLIVA, chose the instruction-tuned pioneer FlanT5 as their foundational LLM component.
-Recently, the LLaMA family, known for its open-source accessibility and impressive performance in NLP tasks, has attracted significant attention.
-The release of LLaMA-2 is expected to continue and possibly foster this trend.
-Numerous VLMs, including LLaVA, MiniGPT-4, InstructBLIP, MiniGPT-v2, and CogVLM, have been developed using language models from the LLaMA family, such as Vicuna, Alpava, and LLaMA-2-Chat, as shown in Table 2.
-Connector
+LLM Components.
+During early VLM development, models such as BLIP-2 and InstructBLIP adopted the instruction-tuned FlanT5 as the LLM component.
+Recently, the open-source, high-performing LLaMA family has drawn significant attention, with LLaMA-2’s release likely to amplify this trend.
+Many VLMs, including LLaVA, MiniGPT-v2, and CogVLM, have adopted language models from the LLaMA family, such as Vicuna, Alpaca, and LLaMA-2-Chat.
+Cross-Modal Connectors.
 BLIP-2 introduced the Q-Former component, which was subsequently adopted by InstructBLIP and BLIVA.
@@
 Models like LLaVA, PaLI-X, MiniGPT-v2, and CogVLM have adopted this design.
-Similarly, Qwen-VL has adopted the linear layer connection approach, but it improves its functionality by adding cross-attention mechanisms.
+Similarly, Qwen-VL has adopted the linear layer connection approach with additional cross-attention mechanisms.
 Comparing MLP with Q-Former, the LLaVA-v1.5 study (Liu et al., 2023a) suggests that the Q-Former may have shortcomings in effectively managing the output length of LLMs.
-Training/Tuning
-The prevalent pre-training methodology for VLMs generally consists of three phases.
-Initially, the fundamental pretraining phase utilizes extensive image-text pairs from the Internet to equip the model with general capabilities for visual-language tasks.
-Subsequently, models undergo multi-task fine-tuning, during which the dataset selection is more carefully curated.
-The final stage is centered on instruction tuning.
-It utilizes a limited amount of instructional or conversational data to optimize the model for interactive user engagement.
+Training Methods.
+VLM training typically goes through three phases.
+Initially, the pretraining phase utilizes extensive image-text pairs from the Internet to equip the model with general multimodal capabilities.
+Then, models undergo multi-task fine-tuning with more carefully curated datasets.
+The final stage centers on instruction tuning with a limited amount of instructional or conversational data.
 During these stages, the training components of the models may differ.
-In the initial stage, the connector may update its parameters.
-However, models typically do not update the LLM parameters at this stage due to their substantial magnitude.
+In the first stage, the connector may update its parameters.
+However, models typically do not update the LLM parameters at this stage due to their substantial scale.
 In contrast, many VLMs update the LLM parameters during the final instruction tuning phase.
-Models with frozen LLM include MiniGPT-4 and BLIVA.
+Models with frozen LLM include BLIVA.
 Although CogVLM does not update its own LLM parameters, it incorporates an additional module into the LLM and updates this module’s parameters.
@@
 2.3 Embodied Agents
-LLM-based agents, especially embodied agents, have demonstrated rapid and significant emergence across various scenarios (Xi et al., 2023).
-The specialized embodied components enable these agents to interact with environments during problem-solving by perceiving the surroundings and performing actions.
-These actions can either be autonomously carried out by the agent itself or with the aid of additional tools.
-In this study, we focus on embodied agents that perform actions independently, rather than relying on external tools.
-This survey analyzes the use of LLMs/VLMs, categorizing their tasks, environments, and design strategies for transforming them into autonomous agents proficient in task-solving.
-A systematic summary can be found in Table 3.
+LLM-based agents, especially embodied agents, have demonstrated rapid and significant emergence (Xi et al., 2023).
+The embodied components enable them to interact with environments by perceiving the surroundings and performing actions.
+We focus on embodied agents that perform actions independently rather than relying on external tools.
+This survey analyzes embodied agents by the LLM component, tasks, environments, and design strategies, as summarized in Table 3.
 Name Time LLM Task Environment LM Output Code Interface LM Prompting/Tuning
@@
 Regarding code interface, ➀ is code-free (language-only), and ➁ involves programming code.
-In LM operations, ➀ is for prompting, ➁ for tuning.
+In LM operations, ➀ is for prompting, and ➁ for tuning.
 It is observed that most embodied agents employ LM prompting for mid-level plans in natural language, which are then translated into low-level robotic actions.
-LLM
-GPT-3.5 and GPT-4, accessible through API services, are the favored LLMs for developing embodied agents.
-As shown in Table 3, models such as Inner Monologue (Huang et al., 2022b), ProgPrompt (Singh et al., 2023), LLM-Planner (Song et al., 2022), Matchat (Zhao et al., 2023b), Grounded Decoding (Huang et al., 2023), Voyager (Wang et al., 2023a), GITM (Zhu et al., 2023), and ViLA (Hu et al., 2023) utilize GPT models as the cognitive core for agents’ thinking and planning processes.
-Additionally, ’Code as Policy’ (Liang et al., 2023) employs CodeX, a derivative of the GPT series focusing on programming.
+LLM Components.
+GPT-3.5 and GPT-4, known for advanced cognitive abilities, are the favored LLMs for developing embodied agents.
+Models such as Inner Monologue (Huang et al., 2022b), ProgPrompt (Singh et al., 2023), LLM-Planner (Song et al., 2022), Matchat (Zhao et al., 2023b), Grounded Decoding (Huang et al., 2023), Voyager (Wang et al., 2023a), GITM (Zhu et al., 2023), and ViLA (Hu et al., 2023) utilize GPT models as the cognitive core for thinking and planning.
+Additionally, ‘Code as Policy’ (Liang et al., 2023) employs CodeX, a derivative of the GPT series focusing on programming.
 Another prominent LLM is PaLM, which has been chosen by agents like SayCan (Ichter et al., 2022), Grounded Decoding, and RT-2 (Brohan et al., 2023).
-Environment and Task
+Environments and Tasks.
 There are two types of environments in which embodied agents operate: simulated and real-world environments.
@@
 Furthermore, numerious embodied agents are able to execute in the physical world.
-These include SayCan, ’Code as Policy’, Inner Monologue, ProgPrompt, Matcha, PaLM-E, Grounded Decoding, RT-2, ViLA, AutoRT (Ahn et al., 2024), and so on.
-From LLM to Agent
+These include SayCan, ‘Code as Policy’, Inner Monologue, ProgPrompt, Matcha, PaLM-E, Grounded Decoding, RT-2, ViLA, AutoRT (Ahn et al., 2024), and so on.
+LLM-based Agents.
 To transform LLMs into agents, a typical method is a two-tier framework: the LLM initially generates natural language sequences as high-level plans, later translated into executable low-level actions.
-Agents using this approach include Inner Monologue, LLM-Planner, Matcha, PaLM-E, Grounded Decoding, ViLA and AutoRT.
+Agents using this approach include Inner Monologue, LLM-Planner, Matcha, PaLM-E, Grounded Decoding, ViLA, and AutoRT.
 Alternatively, some models have implemented an end-to-end action generation scheme.
-Such agents include SayCan, ’Code as Policy’, ProgPrompt, Voyager, GITM, and RT-2.
-Notably, RT-2 treats action as a modality and has introduced the concept of a visual-language-action model.
+Such agents include SayCan, ‘Code as Policy’, ProgPrompt, Voyager, GITM, and RT-2.
+Particularly, RT-2 treats action as a modality and has introduced the concept of a visual-language-action model.
 In addition, while many agents leverage the prompt approach in utilizing LLMs, RT-2 fine-tunes the LLMs for improved performance.
 Moreover, as next-step action generation is inherently logical, some research conveys a preference for using code instead of natural language to facilitate this process.
-Agents that have implemented programming code for action generation include ProgPrompt, ’Code as Policy’, and Voyager.
+Agents that have implemented programming code for action generation include ProgPrompt, ‘Code as Policy’, and Voyager.
 3 Benchmark Evaluation
-In order to explore the question ’Can LMs transform the landscape of VIA?’ and to develop a roadmap for VIALM, we conducted a benchmark evaluation.
+In order to explore the question ‘Can LMs transform the landscape of VIA?’ and to develop a roadmap for VIALM, we conducted a benchmark evaluation.
 This assessment involves measuring and comparing the zero-shot performance of LMs, aiming to help readers understand the models’ strengths and weaknesses in the following four key aspects.
-(1) Viability: Do LMs possess the potential and capabilities necessary for VIA?
-(2) Environment Adaptability: If LMs are effective for VIA, can they maintain their effectiveness in complex environments?
-(3) Functionality: Can LMs achieve effective performance in real-world practice?
-(4) Model Selection: How do different designs of LMs affect performance in VIA?
-Evaluation Models
+(RQ1) Fundamental Skills: Do LMs possess the fundamental multimodal capabilities necessary for VIA?
+(RQ2) Practical Skills: Can LMs achieve effective performance in practice with complex physical environments?
+(RQ3) Model Selection: How do different designs of LMs affect performance in VIA?
+Models and Setup.
 Chen et al.
-(2023b) demonstrated that end-to-end VLMs outperform separate visual and language pipelines in embodied decision-making tasks.
-We selected six SOTA end-to-end VLMs for our evaluation, namely GPT-4, CogVLM, MiniGPT, Qwen-VL, LLaVA and BLIVA.
-The specific versions used were GPT-4, CogVLM-17B-Chat-V1.1, MiniGPT-v2 with LLaMA-2-Chat-7B, Qwen-VL-Chat with Qwen-7B, LLaVA-v1.5-7B, and Bliva-vicuna-7B.
+(2023b) demonstrated that end-to-end VLMs outperform separate visual and language pipelines in embodied tasks.
+We selected six SOTA end-to-end VLMs for our evaluation: GPT-4, CogVLM, MiniGPT, Qwen-VL, LLaVA and BLIVA.
+The specific versions were GPT-4, CogVLM-17B-Chat-V1.1, MiniGPT-v2 with LLaMA-2-Chat-7B, Qwen-VL-Chat with Qwen-7B, LLaVA-v1.5-7B, and Bliva-vicuna-7B.
 All models operated with float16 precision, with the exception of GPT-4.
+We explicitly constructed the prompts, requesting LMs to act as a VI assistant and provide grounded guidance to follow with tactile clues.
 3.1 Evaluation Dataset
@@
 This dataset is designed to (1) accurately reflect, measure, and compare performance on the VIALM task, and (2) remain unexposed to these VLMs during their training and tuning phases.
-Dataset Annotation
+Dataset Annotation.
 Figure 3: Evaluation Data Sample.The test dataset is in the format of a VQA dataset.
@@
 The output answer offers detailed guidance related to this environment to assist the user in achieving this objective.
-We collected images from the Internet of the two most common daily life environments: indoor homes and supermarkets.
-The evaluation dataset was subsequently manually crafted by employing highly educated annotators with English-medium undergraduate or graduate degrees (Li et al., 2023b).
-Figure 3 illustrates a test data sample.
-Dataset Statistics
+We collected images from the Internet of the two most common daily life environments: homes and supermarkets.
+The dataset was subsequently manually curated by employing well-educated and English-fluent annotators, following the practice in Li et al.
+(2023b).
+Specifically, the ground truth construction is supervised by VIA experts at The Hong Kong Society for the Blind (HKSB).
+Dataset Characteristics.
+Each image in the dataset is paired with a corresponding question and an answer.
+Questions are crafted to encourage the models to offer guidance (answers) tailored for VI users.
+In formulating answers, tactile sensation is considered, as previous work indicates that tactile perception is more pronounced among the VI group, compensating for vision loss during task completion (Ottink et al., 2022).
+Figure 3 shows an example from the dataset.
+Dataset Statistics.
 The evaluation dataset consists of 200 samples in total, divided into two scenarios with 100 images each, similar in size to Chen et al.
 (2023b).
-Every image is paired with a corresponding question and answer.
-The questions are designed to prompt the models to offer guidance for VI users, utilizing tactile clues to navigate towards specific targets.
-Notably, the dataset features a broad coverage, with 96 different target objects, such as ’meat’ and ’basin’ in the supermarket scenario, and ’sofa’ and ’fridge’ in the home scenario.
+It features a broad coverage, with 96 different target objects, such as ‘meat’ and ‘basin’ in the supermarket scenario, and ‘sofa’ and ‘fridge’ in the home scenario.
 Furthermore, in comparison to benchmark datasets like OK-VQA (average answer length: 8.1 tokens) (Marino et al., 2019) and CLEVR (average answer length: 18.4 tokens) (Johnson et al., 2017), our dataset exhibits longer questions (average question length: 51 tokens) and significantly longer answers (average answer length: 99.8 tokens).
-It indicates a higher level of difficulty for models in question answering.
+It indicates a higher level of difficulty for models.
 3.2 Evaluation Metrics
 We utilize a hierarchical set of metrics across various dimensions for evaluation. These metrics fall into two main categories: Automatic Evaluation and Human Evaluation.
-Automatic Evaluation Metrics
+Automatic Evaluation Metrics.
 We analyze the generated responses (Ni’mah et al., 2023) using both surface-level and semantic-level metrics.
@@
 For this, we use BERTScore (Zhang et al., 2020), a widely recognized metric that evaluates semantic similarity.
-Human Evaluation Metrics
-The human evaluation measures the models’ performance in three main aspects: Correctness, Actionability, and Clarity, culminating in an overall helpfulness score.
-Correctness evaluates the accuracy of the guidance in describing the environment and indicating target positions.
-Actionability evaluates the user’s ease in following the provided guidance.
-Clarity assesses the comprehensibility of the answer’s language from the user’s perspective.
-The overall helpfulness score is calculated using weights of 0.5 for Actionability, 0.4 for Correctness, and 0.1 for Clarity.
+Human Evaluation Metrics.
+[…]
 4 Experiment Results
 Figure 4: Evaluation Results.
-(a) Overall average experimental results for six models—GPT-4, CogVLM, MiniGPT, Qwen-VL, LLaVA, BLIVA—based on ROUGE, BERTScore, and Human Evaluation metrics.
-(b) Human evaluation results of the six models focusing on the aspects of Correctness, Actionability, and Clarity.
-We present the overall experimental results in Figure 4 (a).
-First, GPT-4 and LLaVA both exhibit relatively high BERTSCore values (above 0.6 within a range of 0 to 1), indicating that current SOTA VLMs possess the potential for the VIALM task (Viability).
-Second, regarding ROUGE, all models demonstrate similar but low values (below 0.2 within a range of 0 to 1), suggesting that the model predictions may be further improved.
-Third, the human evaluation results, which have been scaled into a range from 0 to 1, reveal that GPT-4 performs the best, with LLaVA and CogVLM following as the second-best performers, motivating further study.
+(a) The overall average automatic evaluation results for these six models: GPT-4, CogVLM, MiniGPT, Qwen-VL, LLaVA, and BLIVA, based on ROUGE and BERTScore metrics.
+(b) The human evaluation results of the six models focusing on the aspects of Correctness, Actionability, and Clarity.
 4.1 Automatic Evaluation Results
@@
 The highest and runner-up metric values are distinctly marked in gray.
-’RG’ represents ROUGE, ’BERT_S’ stands for BERTScore, and ’Len’ signifies output length.
-Table 4 displays automatic evaluation results for the two distinct scenarios.
-In both situations,GPT-4 and LLaVA possess the highest BERTScores,indicating their superior abilities.
-Additionally, when considering ROUGE, we can see: (1) all models perform better in the home scenario setting; (2) LLaVA displays relatively high ROUGE values and CogVLM excels in a more challenging environment; (3) GPT-4 performs worse than smaller-scaled models.
-The first observation suggests that the increased visual complexity of the supermarket environment (i.e., a greater variety of objects and more colors in the image) makes the VIALM task more challenging, highlighting the need for models to generalize with environment adaptability.
-Combining the second and third observations and analyzing the generated guidance, we identify GPT-4’s significant drawback: its output is excessively redundant and lengthy, resulting in low ROUGE scores and reduced functionality.
+‘RG’ represents ROUGE, ‘BERT_S’ stands for BERTScore, and ‘Len’ signifies output length.
+The overall automatic experimental results are illustrated in Figure 4 (a).
+We observe GPT-4 and LLaVA exhibit relatively high BERTSCore values (above 0.6 within a range of 0 to 1).
+For ROUGE, all models demonstrate consistently low values (below 0.2 within a range of 0 to 1).
+Given that BERTScore quantifies semantic similarity while ROUGE measures strict matching to the references, this observation illustrates that LMs can generally address the task, though without strict matching, indicating LMs’ fundamental VIA skills (RQ1).
+In a detailed analysis, Table 4 presents automatic evaluation results for the two distinct scenarios.
+In both situations, GPT-4 and LLaVA possess the highest BERTScores, demonstrating superior semantic capabilities compared to other models.
+Additionally, regarding ROUGE, we can see: (1) all models perform better in the home scenario setting; (2) LLaVA displays relatively high ROUGE values and CogVLM excels in a more challenging environment; (3) GPT-4 performs worse than smaller-scaled models.
+The first observation suggests that the increased visual complexity of the supermarket environment (i.e., a greater variety of objects and more colors in the image) makes the VIALM task more challenging, highlighting the need for models to generalize with in more challenging environments (RQ2).
+Combining the second and third observations and analyzing the generated guidance, we identify GPT-4’s significant drawback: its output is excessively redundant and lengthy, resulting in low ROUGE scores.
 In contrast, CogVLM generates the shortest outputs, while LLaVA exhibits the best performance with a moderate guidance length and high ROUGE scores.
-4.2 Human Evaluation Results
 Model Not Grounded Rate Not Fine-grained Rate
@@
 Table 5: Human Evaluation Fail Rates.
-’Not Grounded Rate’ is calculated by the sample proportion with Correctness $\leq$ 3 (Score range: 1–5).
-’Not Fine-grained Rate’ is derived from Actionability $\leq$ 3.
+‘Not Grounded Rate’ is calculated by the sample proportion with Correctness $\leq$ 3 (Score range: 1–5).
+‘Not Fine-grained Rate’ is derived from Actionability $\leq$ 3.
+4.2 Human Evaluation Results
 Figure 4 (b) presents the human evaluation results.
 To mitigate bias, two independent participants conducted evaluations.
-In each distinct scenario, 50 random samples were selected.
+In each scenario, 50 random samples were selected.
 Pearson and Spearman correlation coefficients between the two annotators for each dimension range from 0.7 to 0.9, indicating a strong consensus on the models’ performance.
@@
 Second, as shown in Figure 4 (b), GPT-4 excels in performance across all three dimensions, with CogVLM and LLaVA closely following in second place.
-Subsequently, we analysis the results for the three specific dimensions:
+Subsequently, we analyze the results for the three specific dimensions:
 $\bullet$ In Correctness, GPT-4 performs the best, with CogVLM slightly surpassing LLaVA.
-It might indicate that CogVLM understands the visual environment better than LLaVA, and therefore leads to a superior environment grounding ability.
-$\bullet$ In Actionability, GPT-4 performs the best, with LLaVA narrowly outperforming CogVLM.
-Since we prompted the use of tactile clues, this observation implies LLaVA’s LLM component, LLaMA-2, generates more effective guidance to VI users, incorporating tactile and fine-grained instructions.
+It indicates that CogVLM understands the visual environment better than LLaVA, and therefore leads to a superior environment grounding ability.
+[…]
 $\bullet$ In Clarity, GPT-4 obtains the highest score.
@@
 In contrast, other models provide responses that are not excessively long but tend to be overly simplistic, lacking sufficient information.
-This analysis reveals two prevalent shortcomings shared by all six models, summarized as ’not grounded’(indicating low correctness) and ’not fine-grained’ (indicating low actionability).
-The failure rates are presented in Table 5.
+This analysis reveals two limitations shared by all models, summarized as the answer to RQ2: ‘not grounded’ (indicated by low correctness) and ‘not fine-grained’ (indicated by low actionability).
+Table 5 presents the failure rates.
 It is notable that, despite GPT-4’s superior performance, 25.7% of its responses are not grounded and 32.1% not fine-grained.
-4.3 Case Study
-To demonstrate the models’ output, we present an example case in Figure 5 where the question seeks guidance to obtain mustard.
-The predictions of GPT-4, CogVLM, and LLaVA are provided, yielding the following observations: (1) GPT-4 generates numerous tokens offering general guidance but lacks image grounding.
-In addition, its guidance is not easy for VI users to follow.
-(2) CogVLM exhibits better environment grounding, specifying the object’s position to the user’s left, yet it lacks detailed step-by-step guidance.
-(3) LLaVA’s output emphasizes guiding the user to the specific object but appears to have limited understanding of the environment.
+Additionally, only GPT-4 with a very large scale is capable of providing some general tactile descriptions, while the other five smaller models largely lack this ability.
+Case Study.
 Figure 5: An example of predictions from the top three models.
@@
 CogVLM exhibits advantages in image grounding, while LLaVA emphasizes providing step-by-step guidance for VI users.
+To better understand the above results, we present an example case in Figure 5 where the question seeks guidance to obtain mustard at a supermarket.
+The predictions of GPT-4, CogVLM, and LLaVA yield the following observations: (1) GPT-4 generates long output with general guidance but lacks image grounding.
+(2) CogVLM exhibits better environment grounding, correctly specifying the object’s position, yet it lacks detailed step-by-step guidance.
+(3) LLaVA’s output emphasizes guiding the user to the specific object but appears to have limited understanding of the environment.
+(4) GPT-4 generates general tactile guidance, whereas CogVLM and LLaVA entirely ignore tactile descriptions.
 4.3 Discussions
+Grounded Fine-grained
+Model RG_L BERT_S RG_L BERT_S
+Supermarket
+GPT-4 0.072 0.245 0.115 0.588
+CogVLM 0.144 0.14 0.172 0.495
+MiniGPT 0.102 0.202 0.136 0.485
+Qwen-VL 0.09 0.169 0.138 0.495
+LLaVA 0.097 0.26 0.145 0.554
+BLIVA 0.062 0.213 0.107 0.463
 Home
-Model Rouge_1 Rouge_2 Rouge_L BERTScore
-Grounded Part
-GPT-4 0.156 0.028 0.104 0.405
-CogVLM 0.262 0.043 0.158 0.284
-MiniGPT4 0.211 0.053 0.139 0.415
-Qwen-VL 0.201 0.028 0.132 0.311
-LLaVA 0.186 0.039 0.123 0.404
-BLIVA 0.133 0.025 0.091 0.375
-Fine-Grained Part
-GPT-4 0.292 0.072 0.166 0.573
-CogVLM 0.317 0.083 0.189 0.443
-MiniGPT4 0.295 0.062 0.177 0.455
-Qwen-VL 0.331 0.088 0.193 0.506
-LLaVA 0.341 0.094 0.202 0.544
-BLIVA 0.259 0.062 0.159 0.514
+GPT-4 0.104 0.405 0.166 0.573
+CogVLM 0.158 0.284 0.189 0.443
+MiniGPT 0.139 0.415 0.177 0.455
+Qwen-VL 0.132 0.311 0.193 0.506
+LLaVA 0.123 0.404 0.202 0.544
+BLIVA 0.091 0.375 0.159 0.514
 Table 6: Automatic Evaluation Results with separated guidance parts extracted from ground truths. Metric scores are calculated by comparing model predictions with each corresponding part.
-This section explores model selection strategies to improve VIALM, particularly in addressing the two mentioned limitations.
-We divided the ground truth text into two parts: the grounded part (i.e., environmental), and the fine-grained part (i.e., tactile).
-We then compared model predictions with each part and calculated automatic metric scores.
-Since Table 4 reveals that Qwen-VL and LLaVA exhibit superior performance under the home scenario, a notable deviation from the supermarket scenario and our human evaluation results, we focus on the home scenario.
-Results are displayed in Table 6.
-Analyzing Table 6 reveals two significant findings.
-Firstly, CogVLM excels in environment grounding, possibly due to two factors: (1) the incorporation of the more recent Eva-Clip ViT model, enhancing its visual perception capabilities, and (2) the inclusion of a Visual Expert module in CogVLM’s LLM architecture, augmenting its visual-language alignment abilities.
-Secondly, LLaVA demonstrates superior efficiency in generating fine-grained guidance, possibly because its LLM component, LLaMA-2, might possess advanced capabilities concerning tactile description.
-These initial findings can assist researchers in developing VIALM models.
+[…]
+It is observed that CogVLM performs well in environment grounding and LLaVA excels in generating fine-grained outputs.
+CogVLM’s success in environmental grounding might stem from two factors: (1) the incorporation of the more recent Eva-Clip ViT model, enhancing visual perception, and (2) the inclusion of a Visual Expert module in CogVLM’s LLM architecture, augmenting visual-language alignment.
+LLAVA’s effectiveness in producing fine-grained guidance might come from its advanced LLM component, LLaMA-2.
+Enhancing environment grounding demands advanced visual capabilities, and generating fine-grained guidance necessitates better language abilities.
+Existing models are limited in synergizing both vision and language skills.
+It requires future efforts to develop LMs with environment-grounded multimodal abilities for providing fine-grained guidance in VIA.
 5 Conclusion
-In this study, we defined the VIALM task and carried out a comprehensive survey of LMs for VIA, encompassing LLMS, VLMs, and embodied agents.
+In this study, we defined the task of VIALM and carried out a comprehensive survey of LMs for VIA, encompassing LLMS, VLMs, and embodied agents.
 Furthermore, we constructed a benchmark dataset with comprehensive coverage and high difficulty.
-Subsequently, we conducted a benchmark evaluation to measure the zero-shot performance of LMs, aiming to examine the potential and limitations of current SOTA models in the VIALM task.
+Following this, we conducted a benchmark evaluation to measure the zero-shot performance of LMs, aiming to examine the potential and limitations of current SOTA models in the VIALM task.
 The experimental results revealed two primary limitations in the models: their ability to generate grounded and fine-grained guidance.
-To address these limitations, the experiment results suggest that improving visual processing can aid in producing more grounded advice, and incorporating the tactile modality can contribute to generating more fine-grained guidance for VI users.